OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (mertsj):

First find the slope

OpenStudy (anonymous):

m=1/3 ?

OpenStudy (mertsj):

Now choose either point and plug into: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

y-1=1/3(x-5) ?

OpenStudy (mertsj):

Now distribute the 1/3 and then put it in the form ax+by=c

OpenStudy (anonymous):

y-1=1/3x-5/3 ?